# SPHERE: Unveiling Spatial Blind Spots in Vision-Language Models Through Hierarchical Evaluation

### v3

@@ 4.1 Setup
We describe the experimental setup used to evaluate the performance of state-of-the-art VLMs on SPHERE tasks. All experiments were performed on a NVIDIA GeForce RTX 3090 GPU.
Models: Our evaluation encompasses two model categories: general-purpose VLMs and VLMs with enhanced spatial capabilities. The general-purpose VLMs evaluated are Phi-3.5-Vision Abdin et al. (2024), LLaVA-NeXT Liu et al. (2024), LLaVA-OneVision Li et al. (2024a), InstructBLIP Dai et al. (2023), Idefics2 HuggingFaceM4 (2024), InternVL2.5 Chen et al. (2024b), Qwen-VL Bai et al. (2023), Qwen2-VL Wang et al. (2024), Llama-3.2-Vision Meta (2024), Gemini 2.0 Flash Google (2024), and GPT-4o OpenAI (2024). The VLMs specifically trained for spatial understanding and reasoning are SpatialBot (Cai et al., 2024), SpaceMantis AI (2024b), and SpatialRGPT Cheng et al. (2024). For SpatialRGPT, we use the RGB-only version as the model requires highly accurate mask proposals to make use of depth maps. All model implementations follow their released versions on Hugging Face.
@@ 4.2 Results
The proprietary models, GPT-4o and Gemini 2.0 Flash, outperform the open-weights models evaluated in almost all tasks. Interestingly, amongst open-weights models, we observe that the smaller model LLaVA-OneVision (7B) can outperform larger models in terms of overall accuracy, and performs on par with Gemini 2.0 Flash.
@@ 4.3 Further Analysis
"Given that for the question: <intermediate step question> The answer is: <intermediate step answer>. <final step question>"
For the example in Figure 3, the prompt will be "Given that for the question: Which object is thicker? Fire hydrant or tree trunk? The answer is: Tree trunk. Where can a child be hiding? Behind the fire hydrant or behind the tree?"
@@ 7 Ethics Statement
Acknowledgement
We thank Chanwoo Hwang from Hanyang University for valuable discussions.
@@ Appendix B Detailed Evaluation Results
Model Single-skill Multi-skill Reasoning Pos. Count. Dist. Size Avg P + C D + C D + S Avg Occl. Manip. Avg Overall General purpose VLMs Phi-3.5-Vision (4B) 100.0 100.0 99.7 100.0 99.9 100.0 100.0 99.0 99.7 100.0 100.0 100.0 99.9 LLaVA-NeXT (7B) 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.8 99.6 100.0 100.0 100.0 99.9 LLaVA-OneVision (7B) 99.7 100.0 99.8 98.4 99.5 100.0 100.0 97.3 99.1 99.7 100.0 99.8 99.4 Qwen2-VL (7B) 97.4 100.0 98.3 99.1 98.7 100.0 100.0 98.5 99.5 98.5 99.5 99.0 99.0 InstructBLIP (8B) 97.1 100.0 100.0 99.7 99.2 99.4 100.0 99.9 99.8 99.8 100.0 99.9 99.6 Idefics2 (8B) 92.9 93.0 96.5 91.9 93.6 99.4 96.8 91.4 95.9 98.7 96.0 97.3 95.3 InternVL2.5 (8B) 98.2 100.0 99.4 98.9 99.1 100.0 100.0 99.3 99.8 99.8 100.0 99.9 99.5 Qwen-VL (10B) 99.0 100.0 99.5 100.0 99.6 100.0 100.0 98.4 99.5 96.8 98.8 97.8 99.1 Llama-3.2-Vision (11B) 99.0 100.0 99.0 98.3 99.1 100.0 100.0 98.6 99.5 100.0 99.9 99.9 99.4 Qwen2-VL (72B) 97.6 100.0 99.5 99.0 99.0 100.0 100.0 99.3 99.8 99.6 100.0 99.8 99.5 Llama-3.2-Vision (90B) 95.9 100.0 99.5 98.3 98.4 100.0 100.0 98.3 99.4 100.0 98.8 99.4 99.0 Gemini 2.0 Flash 99.4 100.0 100.0 100.0 99.9 100.0 100.0 98.5 99.5 100.0 100.0 100.0 99.8 GPT-4o 99.2 99.5 98.5 97.0 98.5 100.0 100.0 94.0 98.0 98.5 100.0 99.2 98.5 Spatial VLMs SpatialBot-RGB (3B) 99.2 99.0 99.8 98.4 99.1 100.0 100.0 99.0 99.7 100.0 99.5 99.7 99.4 SpatialBot (3B) 99.2 100.0 100.0 100.0 99.8 100.0 99.4 99.3 99.6 100.0 100.0 100.0 99.8 SpaceMantis (8B) 99.4 100.0 100.0 99.7 99.8 100.0 100.0 98.5 99.5 100.0 100.0 100.0 99.7 SpatialRGPT-RGB (8B) 98.9 100.0 100.0 99.1 99.5 100.0 100.0 99.2 99.7 97.9 95.9 96.9 98.9

### v4

@@ 4.1 Setup
We describe the experimental setup used to evaluate the performance of state-of-the-art VLMs on SPHERE tasks. All experiments were performed on NVIDIA GeForce RTX 3090 GPUs.
Models: Our evaluation encompasses two model categories: general-purpose VLMs and VLMs with enhanced spatial capabilities. The general-purpose VLMs evaluated are Phi-3.5-Vision Abdin et al. (2024), LLaVA-NeXT Liu et al. (2024), LLaVA-OneVision Li et al. (2024a), InstructBLIP Dai et al. (2023), Idefics2 HuggingFaceM4 (2024), InternVL2.5 Chen et al. (2024b), Janus-Pro Chen et al. (2025), Qwen-VL Bai et al. (2023), Qwen2-VL Wang et al. (2024), Qwen2.5-VL Bai et al. (2025), Llama-3.2-Vision Meta (2024), Gemini 2.0 Flash Google (2024), and GPT-4o OpenAI (2024). The VLMs specifically trained for spatial understanding and reasoning are SpatialBot (Cai et al., 2024), SpaceMantis AI (2024b), and SpatialRGPT Cheng et al. (2024). For SpatialRGPT, we use the RGB-only version as the model requires highly accurate mask proposals to make use of depth maps. All model implementations follow their released versions on Hugging Face. We use the default prompt templates and hyperparameters in the demos of the respective models.
The text instructions given take the form:
“<question> Answer the question directly.”
For example, for the first image in Figure 4, the text instruction given is “From the perspective of the viewer, where is the girl relative to the man? Left or Right? Answer the question directly.” We added the instruction “Answer the question directly” to reduce the refusal-to-answer rate and to encourage concise answers that align with the options provided.
@@ 4.2 Results
The proprietary model GPT-4o outperforms the open-weights models evaluated in almost all tasks and has an overall accuracy of 67.9%. Qwen2.5-VL (72B) has the second highest overall accuracy of 64.3%. Interestingly, amongst open-weights models, we observe that the smaller model LLaVA-OneVision (7B) can outperform larger models in terms of overall accuracy, and performs on par with Gemini 2.0 Flash.
@@
In general, we note that the spatial VLMs evaluated performed comparably with or better than most similarly-sized general-purpose VLMs, except for LLaVA-OneVision. In particular, SpatialBot (3B) performed comparably with or better than larger models with up to 11B parameters. This shows that the techniques in spatial VLMs can benefit spatial capabilities. Possible reasons on why the spatial VLMs do not excel on the SPHERE tasks may be the lack of training on the specific skills emphasized by the SPHERE tasks, and the lack of model generalization to the SPHERE tasks.
@@ 4.3 Further Analysis
“Given that for the question: <intermediate step question> The answer is: <intermediate step answer>. <final step question> Answer the question directly.”
For the example in Figure 3, the prompt will be “Given that for the question: Which object is thicker? Fire hydrant or tree trunk? The answer is: Tree trunk. Where can a child be hiding? Behind the fire hydrant or behind the tree? Answer the question directly.”
@@ 7 Ethics Statement
Acknowledgements
We thank Chanwoo Hwang from Hanyang University for valuable discussions. This research is partially supported by the Nanyang Associate Professorship and the National Research Foundation Fellowship (NRF-NRFF13-2021-0006), Singapore.
@@ Appendix B Detailed Evaluation Results
Model Single-skill Multi-skill Reasoning Pos. Count. Dist. Size Avg P + C D + C D + S Avg Occl. Manip. Avg Overall General purpose VLMs Phi-3.5-Vision (4B) 100.0 100.0 99.7 100.0 99.9 100.0 100.0 99.0 99.7 100.0 100.0 100.0 99.9 LLaVA-NeXT (7B) 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.8 99.6 100.0 100.0 100.0 99.9 LLaVA-OneVision (7B) 99.7 100.0 99.8 98.4 99.5 100.0 100.0 97.3 99.1 99.7 100.0 99.8 99.4 Qwen2-VL (7B) 97.4 100.0 98.3 99.1 98.7 100.0 100.0 98.5 99.5 98.5 99.5 99.0 99.0 Qwen2.5-VL (7B) 94.0 100.0 98.8 98.8 97.9 100.0 99.4 96.8 98.7 98.4 97.3 97.9 98.2 Janus-Pro (7B) 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.8 99.9 99.8 100.0 99.9 100.0 InstructBLIP (8B) 97.1 100.0 100.0 99.7 99.2 99.4 100.0 99.9 99.8 99.8 100.0 99.9 99.6 Idefics2 (8B) 92.9 93.0 96.5 91.9 93.6 99.4 96.8 91.4 95.9 98.7 96.0 97.3 95.3 InternVL2.5 (8B) 98.2 100.0 99.4 98.9 99.1 100.0 100.0 99.3 99.8 99.8 100.0 99.9 99.5 Qwen-VL (10B) 99.0 100.0 99.5 100.0 99.6 100.0 100.0 98.4 99.5 96.8 98.8 97.8 99.1 Llama-3.2-Vision (11B) 99.0 100.0 99.0 98.3 99.1 100.0 100.0 98.6 99.5 100.0 99.9 99.9 99.4 Qwen2-VL (72B) 97.6 100.0 99.5 99.0 99.0 100.0 100.0 99.3 99.8 99.6 100.0 99.8 99.5 Qwen2.5-VL (72B) 99.2 100.0 100.0 100.0 99.8 100.0 99.2 99.2 99.5 99.6 99.9 99.8 99.7 Llama-3.2-Vision (90B) 95.9 100.0 99.5 98.3 98.4 100.0 100.0 98.3 99.4 100.0 98.8 99.4 99.0 Gemini 2.0 Flash 99.4 100.0 100.0 100.0 99.9 100.0 100.0 98.5 99.5 100.0 100.0 100.0 99.8 GPT-4o 99.2 99.5 98.5 97.0 98.5 100.0 100.0 94.0 98.0 98.5 100.0 99.2 98.5 Spatial VLMs SpatialBot-RGB (3B) 99.2 99.0 99.8 98.4 99.1 100.0 100.0 99.0 99.7 100.0 99.5 99.7 99.4 SpatialBot (3B) 99.2 100.0 100.0 100.0 99.8 100.0 99.4 99.3 99.6 100.0 100.0 100.0 99.8 SpaceMantis (8B) 99.4 100.0 100.0 99.7 99.8 100.0 100.0 98.5 99.5 100.0 100.0 100.0 99.7 SpatialRGPT-RGB (8B) 98.9 100.0 100.0 99.1 99.5 100.0 100.0 99.2 99.7 97.9 95.9 96.9 98.9
Image
Question
Answer
If you are the man in the image, where is the girl relative to your position? Left or Right?
Ground-truth: Left
GPT-4o: Right ✗
Gemini 2.0 Flash: Right ✗
From the perspective of the viewer, where is the girl relative to the man? Left or Right?
Ground-truth: Right
GPT-4o: Right ✓
Gemini 2.0 Flash: Right ✓
If you are the man pouring wine, where is the man with glasses relative to your position? Left or Right?
Ground-truth: Left
GPT-4o: Right ✗
Gemini 2.0 Flash: Right ✗
Image
Question
Answer
How many buses are visible?
Ground-truth: 0
GPT-4o: 0 ✓
Gemini 2.0 Flash: 0 ✓
How many trees are visible?
Ground-truth: 4
GPT-4o: 2 ✗
Gemini 2.0 Flash: 3 ✗
How many crows can be seen?
Ground-truth: 3
GPT-4o: 3 ✓
Gemini 2.0 Flash: 3 ✓
Image
Question
Answer
Which object is more distant from the plant? Chair or Table?
Ground-truth: Chair
GPT-4o: Chair ✓
Gemini 2.0 Flash: Chair ✓
Which group of people is farther away from the motorcycle? People sitting on the left or People standing on the right?
Ground-truth: People sitting on the left
GPT-4o: People standing on the right ✗
Gemini 2.0 Flash: People sitting on the left ✓
Which boat has a shorter distance to the bank? Boat with a cleaning sign or Boat with passengers?
Ground-truth: Boat with passengers
GPT-4o: Boat with a cleaning sign ✗
Gemini 2.0 Flash: Boat with a cleaning sign ✗
Image
Question
Answer
Which computer is the smallest? The leftmost one or The rightmost one?
Ground-truth: The leftmost one
GPT-4o: The leftmost one ✓
Gemini 2.0 Flash: The leftmost one ✓
Is the fish icon on the ground bigger than the star icon in the picture? Yes or No?
Ground-truth: Yes
GPT-4o: Yes ✓
Gemini 2.0 Flash: Yes ✓
Which cup is higher? The left one or The right one?
Ground-truth: The right one
GPT-4o: The left one ✗
Gemini 2.0 Flash: The left one ✗
Image
Question
Answer
From the perspective of the viewer, how many laptops are on the left side of the man?
Ground-truth: 2
GPT-4o: 2 ✓
Gemini 2.0 Flash: 1 ✗
If you are the man in the image, how many laptops are on your left?
Ground-truth: 0
GPT-4o: 2 ✗
Gemini 2.0 Flash: 2 ✗
How many displays are to the right of the keyboard?
Ground-truth: 1
GPT-4o: 2 ✗
Gemini 2.0 Flash: 3 ✗
Image
Question
Answer
From the perspective of the person, how many horses are farther from him than the fence?
Ground-truth: 3
GPT-4o: 3 ✓
Gemini 2.0 Flash: 2 ✗
From the perspective of the person, how many horses are closer to him than the fence?
Ground-truth: 0
GPT-4o: 1 ✗
Gemini 2.0 Flash: 1 ✗
How many crows are on the railing farther from the viewer?
Ground-truth: 1
GPT-4o: 3 ✗
Gemini 2.0 Flash: 2 ✗
Image
Question
Answer
Are the turqoise plates likely the same size? Yes or No?
Ground-truth: Yes
GPT-4o: Yes ✓
Gemini 2.0 Flash: Yes ✓
The chairs on the top half of the picture are differently sized? Yes or No?
Ground-truth: No
GPT-4o: Yes ✗
Gemini 2.0 Flash: Yes ✗
Which pillow is bigger? The left one or The right one or They are similar in size?
Ground-truth: They are similar in size
GPT-4o: The left one ✗
Gemini 2.0 Flash: The right one ✗
Image
Question
Answer
Where can the pepper shaker most likely be found? Behind the ketchup bottle or Behind the margarita glass?
Ground-truth: Behind the ketchup bottle
GPT-4o: Behind the ketchup bottle ✓
Gemini 2.0 Flash: Behind the ketchup bottle ✓
Where can another teddy bear be found? Inside the suitcase on the white chair or Under the blue dress?
Ground-truth: Inside the suitcase on the white chair
GPT-4o: Under the blue dress ✗
Gemini 2.0 Flash: Under the blue dress ✗
Is it likely that more glasses are hidden behind the pizza boxes? Yes or No?
Ground-truth: No
GPT-4o: No ✓
Gemini 2.0 Flash: Yes ✗
Image
Question
Answer
From the perspective of the womain in pink, can the man wearing glasses stand on her left side? Yes or No?
Ground-truth: Yes
GPT-4o: Yes ✓
Gemini 2.0 Flash: Yes ✓
Can the dog lie between the two cats? Yes or No?
Ground-truth: No
GPT-4o: Yes ✗
Gemini 2.0 Flash: No ✓
From the perspective of the viewer, can we pour all the wine from the nearer glass into the one behind? Yes or No?
Ground-truth: Yes
GPT-4o: Yes ✓
Gemini 2.0 Flash: No ✗